# Quantum plasmonics model of refractive index sensing using photon correlations

### v1

@@ IV Photodetection
with the steady-state mean photon number for any given time in the near field of the MNP as (see Supplementary Information, Section 5)
@@
Alternatively, the polarization operator, given in Refs. [27, 28] as $\hat{P}^{+}=\chi\hat{a}+\mu\hat{\sigma}\approx\chi\hat{a}$ (since $\chi\gg\mu$, for example, in our model, $\chi\approx 64\mu$ and $\mu=72$ D), can be used to calculate the scattered intensity, which is proportional to the expectation value of $\hat{P}^{-}\hat{P}^{+}$, where $\hat{P}^{+}$ is the operator that corresponds to the radiative plasmon mode (since $\chi\propto\sqrt{\gamma_{r}}$ via the Wigner–Weisskopf formula [25, 63]—see Supplementary Information, Section 5). However, the scattered flux in Eq. (15) is more useful, since in addition to representing the radiative part, it can be converted to the mean photocount — a dimensionless parameter that allows us to calculate fluctuations in the scattered light via photodetection theory [35].
It should also be mentioned that $\langle\hat{P}^{-}\hat{P}^{+}\rangle(\text{cf. }\langle\hat{b}^{\dagger}\hat{b}\rangle)$ is proportional to the total scattering spectrum, comprising both the coherent, $|\langle\hat{P}^{+}\rangle|^{2}(\text{cf. }|\langle\hat{b}\rangle|^{2})$, and the incoherent, $\langle\hat{P}^{-}\hat{P}^{+}\rangle-|\langle\hat{P}^{+}\rangle|^{2}(\text{cf.}\langle\hat{b}^{\dagger}\hat{b}\rangle-|\langle\hat{b}\rangle|^{2})$, scattering components [27, 59, 61]. For low light intensities, coherent scattering (scattered light with the same frequency, $\omega_{s}$, as the driving frequency, $\omega$) dominates [27], although the light is only coherent to first order as it is monochromatic. At second order, this monochromatic light can be bunched or antibunched [59], as we show shortly. Thus, in the steady state, we can treat the radiating mode associated with the operator $\hat{b}$ as a single frequency mode ($\omega_{s}=\omega$) for each $\omega$ of the driving field considered, and convert Eq. (15) to the mean photocount, $\langle\hat{m}\rangle$, recorded by a photodetector with a photon collection efficiency $\xi$ (representing the solid angle capture of the radiation and detection efficiency), within some integration time $\mathrm{T}_{int}$ (see Supplementary Information, Section 5), to obtain
@@
In the presence of the QD, a transparency region, where the scattered flux approaches zero, appears in the scattering spectrum at 577$\sim$ 577∼ 577 nm, as shown in Fig. 2(a).
@@
This is referred to as the Fano dip [27, 19] and it is due to destructive interference between the broad plasmon scattering spectrum of the MNP and the narrow QD emission lineshape. The inset of Fig. 2(a) shows the asymmetric Fano spectrum, which begins at the Fano dip and reaches a peak at 577.038$\sim$ 577.038∼ 577.038 nm, referred to as the Fano peak [27, 19], where the plasmon spectrum interferes constructively with the QD emission. The driving field intensity, $I_{0}\sim 33.6$ W/cm2, is chosen in our analysis, as it leads to a scattering spectrum in which the ratio of the scattered flux at the LSPR, 535.5$\sim$ 535.5∼ 535.5 nm, to the scattered flux at the Fano resonance, 577.038$\sim$ 577.038∼ 577.038 nm, tends to unity, as shown in Fig. 2(a). This is useful in our plasmonic sensing model since differences in intensity can lead to differences in the relative noise (and thus the signal-to-noise ratio which influences the sensor’s resolution) and we would like to draw an unbiased comparison of the intensity-based sensing at driving wavelengths within the plasmon and Fano spectra. The standard deviation in the mean photocount for a single frequency mode is [35]
@@
is the second factorial moment of the photocount operator in the steady state for small ${\rm T}_{int}$ and the two-photon scattering probability density, $\langle(\hat{b}^{\dagger})^{2}\hat{b}^{2}\rangle$, is obtained as (see Supplementary Information, Section 5)
@@
We now have expressions for $\langle\hat{m}\rangle$, $\Delta m$, and $g^{(2)}(0)$ using Eqs. (17), (18) and (22). The final parameter needed is the noise in $g^{(2)}(0)$, i.e., $\Delta g^{(2)}(0)$, which we discuss soon. First, we study the behaviour of $g^{(2)}(0)$. The photon statistics of the scattered light in terms of $g^{(2)}(\tau)$ is shown in Fig. 2(b), where we have used QuTiP as we only have an analytical formula for $g^{(2)}(0)$, as given in Eq. (22). Studying $g^{(2)}(\tau)$ provides additional information about the temporal behavior of the second-order statistics of the scattered light from the system, although ultimately we are interested in using $g^{(2)}(0)$ only as a sensor signal. Note that due to the time delay, $\tau$, $g^{(2)}(\tau)$ depends on the intermediate dynamics of the MNP-QD system, in which changes in $\hat{\rho}$ can occur on a timescale as small as $\gamma_{pl}^{-1}$ (5.75$\sim$ 5.75∼ 5.75 fs), even though the radiative flux is determined by the lifetime $\gamma_{r}^{-1}$ (4.29$\sim$ 4.29∼ 4.29 ps).
Within the plasmon spectrum, and especially at the LSPR in Fig. 2(a), the scattered light has the same second-order statistics as that of the incident light, i.e., coherent light, and $g^{(2)}(\tau)=1$ at all times, associated with a random arrival of photons at the detector according to Poissonian statistics. At the Fano peak and dip it is necessary to check the full time dependence of $g^{(2)}(\tau)$ for the purposes of using $g^{(2)}(0)$ as a sensor signal, as we must consider how the numerator and the denominator in Eq. (21) are measured in an experiment. Both are measured using photon counting within a finite time period, or integration time, $\mathrm{T}_{int}$, which must be fixed at the same value for the correct normalization. At the Fano peak, $g^{(2)}(\tau)$ varies on the 5105-105 - 10 ps timescale ($\sim\gamma_{r}^{-1}$), as seen in Fig. 2(b), and as we require a stable signal for sensing, the integration time, $\mathrm{T}_{int}=3$ ps, is chosen such that $\mathrm{T}_{int}<\gamma_{r}^{-1}$, within which $g^{(2)}(\tau=0)=g^{(2)}(\tau=\mathrm{T}_{int})\approx 0.17$ (see Supplementary Information, Section 6). From $\tau=0$ to $\tau=\mathrm{T}_{int}$, the scattered light is heavily bunched at the Fano dip, $g^{(2)}(\tau)>1$, but approximately constant, and heavily antibunched at the Fano peak, $g^{(2)}(\tau)\sim 0.17$. At longer delay times, $\tau>\mathrm{T}_{int}$, the second-order statistics of the scattered light at the Fano peak ceases to be uniform as it approaches that of the incident light, while the scattered light at the Fano dip oscillates around $g^{(2)}(\tau)=1$ before second-order coherent scattering finally dominates, in agreement with the behaviour found in Ref. [27].
@@
Here, $g_{D}^{(2)}(0)=g^{(2)}(0)$ if ${\rm T}_{int}$ is small enough such that $g^{(2)}(0)=g^{(2)}(\tau={\rm T}_{int})$, which is satisfied for ${\rm T}_{int}=3$ ps, as discussed above. From Eq. (23), it can be seen that a value of $g^{(2)}(0)<1$ reduces the value of $\Delta m$ below that of the shot noise, $\Delta m=\sqrt{\langle\hat{m}\rangle}$, if the photocount, $\langle\hat{m}\rangle$, is not too small. Using Eq. (23), and ensuring ${\rm T}_{int}$ is the same for all quantities, the photon statistics of the scattered light, as captured by the different values of $g^{(2)}(0)$, can be used to investigate the wavelength dependence of the intensity noise at different regions of the scattering spectrum (from the LSPR to the Fano dip and Fano peak), allowing us to deduce the following: at driving wavelengths within the plasmon spectrum, $\Delta m=\sqrt{\langle\hat{m}\rangle}$, since $g^{(2)}(0)=1$, so that the noise is Poissonian and the same as the shot noise in the driving intensity, while within the Fano spectrum, $\Delta m<\sqrt{\langle\hat{m}\rangle}$, since $g^{(2)}(0)<1$, characterized by a near-regular photocount pattern on the detector, so that the noise is sub-Poissonian and reduced below the shot noise. At the Fano dip, $\Delta m>\sqrt{\langle\hat{m}\rangle}$, since $g^{(2)}(0)>1$, so that the noise is super-Poissonian and thus the refractive index resolution cannot be improved at driving wavelengths within the Fano dip using the intensity as the signal. However, as we will show in Sec. V, besides the driving wavelength, the role of the antibunching effect in reducing the noise also depends strongly on the magnitude of the mean photocount, and thus it is not as straightforward to obtain a reduction in the noise via antibunching as it might at first appear.
Before showing the results of the intensity as a signal and its noise, we complete the derivation of the noise in $g^{(2)}(0)$. When it is used as a sensing signal we have that fluctuations in $g^{(2)}(0)$ will affect the resolution of the sensor. Though $g^{(2)}(0)=\langle\hat{m}(\hat{m}-1)\rangle/\langle\hat{m}\rangle^{2}$ is independent of the detector efficiency, as both $\langle\hat{m}(\hat{m}-1)\rangle$ and $\langle\hat{m}\rangle^{2}$ are proportional to $\xi^{2}$, the standard deviation in $g^{(2)}(0)$ is obtained via the propagation of uncertainties in both $\langle\hat{m}(\hat{m}-1)\rangle$ and $\langle\hat{m}\rangle^{2}$ using the standard error propagation rule in Ref. [68], as (see Supplementary Information, Section 6)
@@
is the standard deviation in the second factorial moment of the photocount, $\langle\hat{m}(\hat{m}-1)\rangle$, and $g^{(3)}(0)$ and $g^{(4)}(0)$ are the third- and fourth-order steady-state correlation functions at zero-delay time, obtained respectively as (see Supplementary Information, Section 5)
@@
The second-, third-, and fourth-order steady-state correlation functions within the Fano spectrum at zero-delay time are shown in Fig. 3. Both $g^{(3)}(0)$ and $g^{(4)}(0)$ show that the scattered light is antibunched at the Fano resonance and at driving wavelengths near the Fano resonance, i.e., $g^{(n)}(0)<1,n=3,4$, in accordance with $g^{(2)}(0)$, where $g^{(4)}(0)<g^{(3)}(0)<g^{(2)}(0)<1$, in agreement with previous works [39].
@@
Figure 3: (Color online) The zero-delay time steady-state correlation functions, $g^{(2)}(0)$, $g^{(3)}(0)$, and $g^{(4)}(0)$, at wavelengths within the Fano spectrum.
@@
respectively, where $\mathcal{N}=1/\mathrm{T}_{int}$ is the number of measurements possible in one second, and for simplicity the role of the detector recovery time in these measurements has been ignored. In an experiment, $\mathrm{T}_{int}$ would be the collection time for $\langle\hat{m}\rangle$ using one single-photon detector and the coincidence window for $g^{(2)}(0)$ using a pair of single-photon detectors. For example, $\langle\hat{m}\rangle\approx 12.42\times 10^{-5}$ in $\mathrm{T}_{int}$ at the LSPR (equivalent to $4.14\times 10^{7}$ photon counts in one second of measurements), as we will show in Sec. V. This corresponds to $\sigma_{m}\approx\langle\hat{m}\rangle^{1/2}(1/{\rm T}_{int})^{-1/2}=(12.42\times 10^{-5})^{1/2}(3\times 10^{-12})^{1/2}\sim 1.93\times 10^{-8}$ in ${\rm T}_{int}$. Thus, the use of repeated measurements reduces the noise $\sigma_{m}$ to below that of the signal $\langle\hat{m}\rangle$.
@@ V Sensing performance
Figure 4: (Color online) Optimal driving wavelengths for intensity sensing at $n=1.3330$ within the (a) plasmon spectrum and (b) Fano spectrum. Upper panels: photocounts, $\langle\hat{m}\rangle$, obtained in $\mathrm{T}_{int}$. Lower panels: second derivative (with respect to $\lambda$) of the photocounts, $\langle\hat{m}\rangle^{\prime\prime}$.
@@
each of which is in units of RIU-1, where RIU stands for refractive index unit. In Eq. (28a), the driving wavelength (equivalent to the scattered wavelength), $\lambda_{D}$, is a wavelength either within the plasmon spectrum or the Fano spectrum, while the driving wavelength in Eq. (28b) is a wavelength within the Fano spectrum only, since the $g^{(2)}(0)$ values within the plasmon spectrum are independent of the refractive index of the medium, i.e., $g^{(2)}(0)=1$ across the plasmon spectrum, as shown in Fig. 2(b). It was reported in Refs. [53, 2, 3] that driving wavelengths corresponding to the inflection points on the plasmon spectrum are the optimal wavelengths for plasmonic sensing. We therefore define the enhancement in the refractive index sensitivity as
@@
In either the plasmon or Fano spectrum, the inflection points correspond to wavelengths where the second derivative (with respect to $\lambda$) of the mean photocount is zero, while the resonance corresponds to the wavelength where the second derivative of the mean photocount reaches a minimum, as indicated in the lower panels of Fig. 4. As we will show shortly, intensity-based sensing for both the plasmon and Fano spectrum is best carried out at driving wavelengths corresponding to the inflection points. Due to the symmetric (plasmon spectrum, Fig. 4(a)) versus asymmetric (Fano spectrum, Fig. 4(b)) lineshapes of the scattering spectrum, the mean photocounts vary at the inflection points. However, the driving intensity has been tuned so that the mean photocount obtained at the LSPR is approximately the same as the value obtained at the middle point of the Fano resonance, i.e., $\langle\hat{m}\rangle\sim 12.42\times 10^{-5}$ counts in $\mathrm{T}_{int}$, in both Figs. 4(a) and 4(b), which allows us to fairly compare the sensing performance at the two regions. On the other hand, the system can be driven such that the plasmon peak is enhanced relative to the Fano peak (increase in $I_{0}$), or such that the plasmon peak is suppressed relative to the Fano peak (decrease in $I_{0}$). Either of these two approaches will lead to an unfair comparison of the sensing performance at the plasmon and Fano resonances since any definition of an enhancement factor will be flawed. More information regarding the dependence of the Fano resonance on the driving field intensity can be found in Refs. [27, 31]. In the top part of Table 2, the wavelengths, $\lambda_{D}$, corresponding to the inflection points have been listed, along with the sensitivities, $S_{I}$, and enhancement factors, $\mathcal{E}$, at these points. Note that the quantum model was essential to obtain the Fano resonance results, and as we will show, to understand whether the photon antibunching effect that occurs within the Fano spectrum can lead to an enhancement in the sensing resolution. The quantum model also opens up the possible use of other observables for sensing, such as $g^{(2)}(0)$, and its associated noise.
@@
Figure 5: (Color online) Driving wavelengths for $g^{(2)}(0)$-based refractive index sensing within the Fano spectrum at $n=1.3330$. Upper panel: $g^{(2)}(0)$. Lower panel: second derivative of $g^{(2)}(0)$ with respect to $\lambda$, $[g^{(2)}(0)]^{\prime\prime}$.
$\lambda_{D}$ (nm) FL =577.031=577.031= 577.031 FM =577.038=577.038= 577.038 FR =577.045=577.045= 577.045
PL =530.770=530.770= 530.770 PM =535.500=535.500= 535.500 PR =540.270=540.270= 540.270
$\mathrm{S}_{I}\times 10^{-4}$FL :4.04:4.04: 4.04 @FM :4.15:4.15: 4.15 @FR :12.17:12.17: 12.17
[RIU -1] @PL :5.87:5.87: 5.87 @PM :3.93:3.93: 3.93 @PR :11.37:11.37: 11.37
[…]
$\mathrm{S}_{I-I}$ [RIU -1] @FL :3.85:3.85: 3.85 @FM :1.05:1.05: 1.05 @FR :0.50:0.50: 0.50
$\Delta n_{I}\times 10^{-5}$FL :4.60:4.60: 4.60 @FM :5.50:5.50: 5.50 @FR :2.10:2.10: 2.10
[RIU] @PL :3.90:3.90: 3.90 @PM :6.90:6.90: 6.90 @PR :2.10:2.10: 2.10
[…]
$\Delta n_{I-I}\times 10^{-3}$ [RIU] @FL :1.3:1.3: 1.3 @FM :2.8:2.8: 2.8 @FR :7.6:7.6: 7.6
@@
Figure 6: (Color online) Intensity sensitivity, standard deviation (noise) for one second of measurements, and sensing resolution (i.e., limit of detection) at driving wavelengths within the plasmon spectrum [(a), (b), (c)], and Fano spectrum [(d), (e), (f)], obtained within the refractive index range, $n=1.3330-1.3334$. The sensing resolution can be improved by taking measurements for more than one second, with the results for the plasmon and Fano resonances scaling by the same factor.

### v2

@@ IV Photodetection
with the steady-state mean photon number for any given time in the near field of the MNP as (see Supplemental Material, Section 5 [57])
@@
Alternatively, the polarization operator, given in Refs. [22, 23] as $\hat{P}^{+}=\chi\hat{a}+\mu\hat{\sigma}\approx\chi\hat{a}$ (since $\chi\gg\mu$, for example, in our model, $\chi\approx 64\mu$ and $\mu=72$ D), can be used to calculate the scattered intensity, which is proportional to the expectation value of $\hat{P}^{-}\hat{P}^{+}$, where $\hat{P}^{+}$ is the operator that corresponds to the radiative plasmon mode (since $\chi\propto\sqrt{\gamma_{r}}$ via the Wigner–Weisskopf formula [20, 58]—see Supplemental Material, Section 5 [57]). However, the scattered flux in Eq. (15) is more useful, since in addition to representing the radiative part, it can be converted to the mean photocount — a dimensionless parameter that allows us to calculate fluctuations in the scattered light via photodetection theory [31].
It should also be mentioned that $\langle\hat{P}^{-}\hat{P}^{+}\rangle(\text{cf. }\langle\hat{b}^{\dagger}\hat{b}\rangle)$ is proportional to the total scattering spectrum, comprising both the coherent, $|\langle\hat{P}^{+}\rangle|^{2}(\text{cf. }|\langle\hat{b}\rangle|^{2})$, and the incoherent, $\langle\hat{P}^{-}\hat{P}^{+}\rangle-|\langle\hat{P}^{+}\rangle|^{2}(\text{cf.}\langle\hat{b}^{\dagger}\hat{b}\rangle-|\langle\hat{b}\rangle|^{2})$, scattering components [22, 51, 53]. For low light intensities, coherent scattering (scattered light with the same frequency, $\omega_{s}$, as the driving frequency, $\omega$) dominates [22], although the light is only coherent to first order as it is monochromatic. At second order, this monochromatic light can be bunched or antibunched [51], as we show shortly. Thus, in the steady state, we can treat the radiating mode associated with the operator $\hat{b}$ as a single frequency mode ($\omega_{s}=\omega$) for each $\omega$ of the driving field considered, and convert Eq. (15) to the mean photocount, $\langle\hat{m}\rangle$, recorded by a photodetector with a photon collection efficiency $\xi$ (representing the solid angle capture of the radiation and detection efficiency), within some integration time $\mathrm{T}_{int}$ (see Supplemental Material, Section 5 [57]), to obtain
@@
In the presence of the QD, a transparency region, where the scattered flux approaches zero, appears in the scattering spectrum at $\sim 577$ nm, as shown in Fig. 2(a).
@@
This is referred to as the Fano dip [22, 30] and it is due to destructive interference between the broad plasmon scattering spectrum of the MNP and the narrow QD emission lineshape. The inset of Fig. 2(a) shows the asymmetric Fano spectrum, which begins at the Fano dip and reaches a peak at $\sim 577.038$ nm, referred to as the Fano peak [22, 30], where the plasmon spectrum interferes constructively with the QD emission. The driving field intensity, $I_{0}\sim 33.6$ W/cm2, is chosen in our analysis, as it leads to a scattering spectrum in which the ratio of the scattered flux at the LSPR, $\sim 535.5$ nm, to the scattered flux at the Fano resonance, $\sim 577.038$ nm, tends to unity, as shown in Fig. 2(a). This is useful in our plasmonic sensing model since differences in intensity can lead to differences in the relative noise (and thus the signal-to-noise ratio which influences the sensor’s resolution) and we would like to draw an unbiased comparison of the intensity-based sensing at driving wavelengths within the plasmon and Fano spectra. The scattered flux of the Fano resonance is large and comparable to the plasmon resonance here mainly due to the size of the dipole moment taken for the QD ($\mu=72$ D) and the driving intensity $I_{0}$ used, which keeps us in the weak-excitation limit.
The standard deviation in the mean photocount for a single frequency mode is [31]
@@
is the second factorial moment of the photocount operator in the steady state for small ${\rm T}_{int}$ and the two-photon scattering probability density, $\langle(\hat{b}^{\dagger})^{2}\hat{b}^{2}\rangle$, is obtained as (see Supplemental Material, Section 5 [57])
@@
We now have expressions for $\langle\hat{m}\rangle$, $\Delta m$, and $g^{(2)}(0)$ using Eqs. (17), (18) and (22). The final parameter needed is the noise in $g^{(2)}(0)$, i.e., $\Delta g^{(2)}(0)$, which we discuss soon. First, we study the behaviour of $g^{(2)}(0)$. The photon statistics of the scattered light in terms of $g^{(2)}(\tau)$ is shown in Fig. 2(b), where we have used QuTiP as we only have an analytical formula for $g^{(2)}(0)$, as given in Eq. (22). Studying $g^{(2)}(\tau)$ provides additional information about the temporal behavior of the second-order statistics of the scattered light from the system, although ultimately we are interested in using $g^{(2)}(0)$ only as a sensor signal. Note that due to the time delay, $\tau$, $g^{(2)}(\tau)$ depends on the intermediate dynamics of the MNP-QD system, in which changes in $\hat{\rho}$ can occur on a timescale as small as $\gamma_{pl}^{-1}$ ($\sim 5.75$ fs), even though the radiative flux is determined by the lifetime $\gamma_{r}^{-1}$ ($\sim 4.29$ ps).
Within the plasmon spectrum, and especially at the LSPR in Fig. 2(a), the scattered light has the same second-order statistics as that of the incident light, i.e., coherent light, and $g^{(2)}(\tau)=1$ at all times, associated with a random arrival of photons at the detector according to Poissonian statistics. At the Fano peak and dip it is necessary to check the full time dependence of $g^{(2)}(\tau)$ for the purposes of using $g^{(2)}(0)$ as a sensor signal, as we must consider how the numerator and the denominator in Eq. (21) are measured in an experiment. Both are measured using photon counting within a finite time period, or integration time, $\mathrm{T}_{int}$, which must be fixed at the same value for the correct normalization. At the Fano peak, $g^{(2)}(\tau)$ varies on the $5-10$ ps timescale ($\sim\gamma_{r}^{-1}$), as seen in Fig. 2(b), and as we require a stable signal for sensing, the integration time, $\mathrm{T}_{int}=3$ ps, is chosen such that $\mathrm{T}_{int}<\gamma_{r}^{-1}$, within which $g^{(2)}(\tau=0)=g^{(2)}(\tau=\mathrm{T}_{int})\approx 0.17$ (see Supplemental Material, Section 6 [57]). From $\tau=0$ to $\tau=\mathrm{T}_{int}$, the scattered light is heavily bunched at the Fano dip, $g^{(2)}(\tau)>1$, but approximately constant, and heavily antibunched at the Fano peak, $g^{(2)}(\tau)\sim 0.17$. At longer delay times, $\tau>\mathrm{T}_{int}$, the second-order statistics of the scattered light at the Fano peak ceases to be uniform as it approaches that of the incident light, while the scattered light at the Fano dip oscillates around $g^{(2)}(\tau)=1$ before second-order coherent scattering finally dominates, in agreement with the behaviour found in Ref. [22].
@@
Here, $g_{D}^{(2)}(0)=g^{(2)}(0)$ if ${\rm T}_{int}$ is small enough such that $g^{(2)}(0)=g^{(2)}(\tau={\rm T}_{int})$, which is satisfied for ${\rm T}_{int}=3$ ps, as discussed above. From Eq. (23), it can be seen that a value of $g^{(2)}(0)<1$ reduces the value of $\Delta m$ below that of the shot noise, $\Delta m=\sqrt{\langle\hat{m}\rangle}$, if the photocount, $\langle\hat{m}\rangle$, is not too small. Using Eq. (23), and ensuring ${\rm T}_{int}$ is the same for all quantities, the photon statistics of the scattered light, as captured by the different values of $g^{(2)}(0)$, can be used to investigate the wavelength dependence of the intensity noise at different regions of the scattering spectrum (from the LSPR to the Fano dip and Fano peak), allowing us to deduce the following: at driving wavelengths within the plasmon spectrum, $\Delta m=\sqrt{\langle\hat{m}\rangle}$, since $g^{(2)}(0)=1$, so that the noise is Poissonian and the same as the shot noise in the driving intensity, while within the Fano spectrum, $\Delta m<\sqrt{\langle\hat{m}\rangle}$, since $g^{(2)}(0)<1$, characterized by a near-regular photocount pattern on the detector, so that the noise is sub-Poissonian and reduced below the shot noise. At the Fano dip, $\Delta m>\sqrt{\langle\hat{m}\rangle}$, since $g^{(2)}(0)>1$, so that the noise is super-Poissonian and thus the refractive index resolution cannot be improved at driving wavelengths within the Fano dip using the intensity as the signal.
Before showing the results of the intensity as a signal and its noise, we complete the derivation of the noise in $g^{(2)}(0)$. When it is used as a sensing signal we have that fluctuations in $g^{(2)}(0)$ will affect the resolution of the sensor. Though $g^{(2)}(0)=\langle\hat{m}(\hat{m}-1)\rangle/\langle\hat{m}\rangle^{2}$ is independent of the detector efficiency, as both $\langle\hat{m}(\hat{m}-1)\rangle$ and $\langle\hat{m}\rangle^{2}$ are proportional to $\xi^{2}$, the standard deviation in $g^{(2)}(0)$ is obtained via the propagation of uncertainties in both $\langle\hat{m}(\hat{m}-1)\rangle$ and $\langle\hat{m}\rangle^{2}$ using the standard error propagation rule in Ref. [64], as (see Supplemental Material, Section 6 [57])
@@
is the standard deviation in the second factorial moment of the photocount, $\langle\hat{m}(\hat{m}-1)\rangle$, and $g^{(3)}(0)$ and $g^{(4)}(0)$ are the third- and fourth-order steady-state correlation functions at zero-delay time, obtained respectively as (see Supplemental Material, Section 5 [57])
@@
The second-, third-, and fourth-order steady-state correlation functions within the Fano spectrum at zero-delay time are shown in Fig. 3. Both $g^{(3)}(0)$ and $g^{(4)}(0)$ show that the scattered light is antibunched at the Fano resonance and at driving wavelengths near the Fano resonance, i.e., $g^{(n)}(0)<1,n=3,4$, in accordance with $g^{(2)}(0)$, where $g^{(4)}(0)<g^{(3)}(0)<g^{(2)}(0)<1$, in agreement with previous works [65]. Note that the narrow wavelength range in Fig. 3 is achievable with the current state of comb laser technology [66].
@@
Figure 3: (Color online) The zero-delay time steady-state correlation functions, $g^{(2)}(0)$, $g^{(3)}(0)$, and $g^{(4)}(0)$, at wavelengths within the Fano spectrum. Here, the Fano spectrum is an enlarged version of the inset in Fig. 2(a) showing only the Fano peak.
@@
respectively, where $\mathcal{N}=1/\mathrm{T}_{int}$ is the number of measurements possible in one second, and for simplicity the role of the detector recovery time in these measurements has been ignored. In an experiment, $\mathrm{T}_{int}$ would be the collection time for $\langle\hat{m}\rangle$ using one single-photon detector and the coincidence window for $g^{(2)}(0)$ using a pair of single-photon detectors. For example, $\langle\hat{m}\rangle\approx 12.42\times 10^{-5}$ in $\mathrm{T}_{int}$ at the LSPR (equivalent to $4.14\times 10^{7}$ photon counts in one second of measurements), as we show in Sec. V. This corresponds to $\sigma_{m}\approx\langle\hat{m}\rangle^{1/2}(1/{\rm T}_{int})^{-1/2}=(12.42\times 10^{-5})^{1/2}(3\times 10^{-12})^{1/2}\sim 1.93\times 10^{-8}$ in ${\rm T}_{int}$. Thus, the use of repeated measurements reduces the noise $\sigma_{m}$ to below that of the signal $\langle\hat{m}\rangle$.
@@ V Sensing performance
Figure 4: (Color online) Optimal driving wavelengths for intensity sensing at $n=1.3330$ within the (a) plasmon spectrum and (b) Fano spectrum showing the Fano peak only. Upper panels: photocounts, $\langle\hat{m}\rangle$, obtained in $\mathrm{T}_{int}$. Lower panels: second derivative (with respect to $\lambda$) of the photocounts, $\langle\hat{m}\rangle^{\prime\prime}$.
@@
each of which is in units of RIU-1, where RIU stands for refractive index unit. In Eq. (28a), the driving wavelength (equivalent to the scattered wavelength), $\lambda_{D}$, is a wavelength either within the plasmon spectrum or the Fano spectrum in Fig. 3, while the driving wavelength in Eq. (28b) is a wavelength within the Fano spectrum in Fig. 3, since the $g^{(2)}(0)$ values within the plasmon spectrum are independent of the refractive index of the medium, i.e., $g^{(2)}(0)=1$ across the plasmon spectrum, as shown in Fig. 2(b). It was reported in Refs. [41, 2, 3] that driving wavelengths corresponding to the inflection points on the plasmon spectrum are the optimal wavelengths for plasmonic sensing. Mathematically, they are points where the second derivative of the intensity with respect to the wavelength is zero. In other words, they correspond to the steepest points on the intensity curve. At these fixed wavelength points, as the plasmon peak shifts left or right with the refractive index, the intensity changes maximally. We therefore define the enhancement in the refractive index sensitivity as
@@
In either the plasmon or Fano spectrum, the inflection points correspond to wavelengths where the second derivative (with respect to $\lambda$) of the mean photocount is zero, while the resonance corresponds to the wavelength where the second derivative of the mean photocount reaches a minimum, as indicated in the lower panels of Fig. 4. Due to the symmetric (plasmon spectrum, Fig. 4(a)) versus asymmetric (Fano spectrum, Fig. 4(b)) lineshapes of the scattering spectrum, the mean photocounts vary at the inflection points. However, the driving intensity has been tuned so that the mean photocount obtained at the LSPR is approximately the same as the value obtained at the middle point of the Fano resonance, i.e., $\langle\hat{m}\rangle\sim 12.42\times 10^{-5}$ counts in $\mathrm{T}_{int}$, in both Figs. 4(a) and 4(b), which allows us to fairly compare the sensing performance at the two regions. On the other hand, the system can be driven such that the plasmon peak is enhanced relative to the Fano peak (increase in $I_{0}$), or such that the plasmon peak is suppressed relative to the Fano peak (decrease in $I_{0}$). Either of these two approaches will lead to an unfair comparison of the sensing performance at the plasmon and Fano resonances since any definition of an enhancement factor will be flawed. More information regarding the dependence of the Fano resonance on the driving field intensity can be found in Refs. [22, 26]. In the top part of Table 2, the wavelengths, $\lambda_{D}$, corresponding to the inflection points have been listed, along with the sensitivities, $S_{I}$, and enhancement factors, $\mathcal{E}$, at these points.
@@
Figure 5: (Color online) Driving wavelengths for $g^{(2)}(0)$-based refractive index sensing within the Fano spectrum (wavelengths around the Fano peak only) at $n=1.3330$. Upper panel: $g^{(2)}(0)$. Lower panel: second derivative of $g^{(2)}(0)$ with respect to $\lambda$, $g^{(2)}(0)^{\prime\prime}$.
$\lambda_{D}$ (nm) FL $=577.031$ FM $=577.038$ FR $=577.045$
PL $=530.770$ PM $=535.500$ PR $=540.270$
$\mathrm{S}_{I}\times 10^{-4}$FL $:4.04$FM $:4.15$FR $:12.17$
[RIU -1] @PL $:5.87$PM $:3.93$PR $:11.37$
$\mathcal{E}_{S}$ $\sim 0.69$ $\sim 1.06$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \sim 1.07$
$\mathrm{S}_{I-I}$ [RIU -1] @FL $:3.85$FM $:1.05$FR $:0.50$
$\Delta n_{I}\times 10^{-5}$FL $:4.60$FM $:5.50$FR $:2.10$
[RIU] @PL $:3.90$PM $:6.90$PR $:2.10$
$\mathcal{E}_{\Delta n}$ $\sim 0.84$ $\sim 1.25$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \sim 1.00$
$\Delta n_{I-I}\times 10^{-3}$ [RIU] @FL $:1.3$FM $:2.8$FR $:7.6$
@@
Figure 6: (Color online) Intensity sensitivity, standard deviation (noise) for one second of measurements, and sensing resolution (i.e., limit of detection) at driving wavelengths within the plasmon spectrum [(a), (b), (c)], and Fano spectrum (wavelengths around the Fano peak only) [(d), (e), (f)], obtained within the refractive index range, $n=1.3330-1.3334$. The sensing resolution can be improved by taking measurements for more than one second, with the results for the plasmon and Fano resonances scaling by the same factor.
@@
These resolutions are obtained by using a simple linear error propagation formula [3]. However, a more rigorous approach is to calculate the Fisher information and obtain the Cramér-Rao bound, which is a lower bound on the estimation error, or resolution using our terminology. The Fisher information for an unknown parameter $\theta$ measured using observable $X$ with a Gaussian probability distribution is $I(\theta)=\Delta^{-2}$, where $\Delta$ is the standard deviation of the probability distribution for $X$ [67, 68]. For ${\cal N}$ independent and identically distributed measurements we then have $I_{\cal N}(\theta)={\cal N}I(\theta)$. The Cramér-Rao bound is then $[I_{\cal N}(\theta)]^{-1/2}=\Delta/\sqrt{\cal N}$. Taking the sensitivities $S_{I}$ and $S_{I-I}$ to be linear for small changes in the refractive index (see Supplemental Material, Section 5 [57]) and the probability distribution for $\langle\hat{m}\rangle$ and $g^{(2)}(0)$ to be Gaussian with standard deviations $\Delta m$ and $\Delta g^{(2)}(0)$, we have that the Cramér-Rao bounds for the resulting refractive index measurements are equal to Eqs. (30a) and (30b), respectively. The sensitivities $S_{I}$ and $S_{I-I}$ therefore allow the propagation of the bounds on the resolution of the photocount and second-order correlation to bounds on the resolution of the refractive index.